OpenStudy (anonymous):

Solve -2(x-3)is greater than or equal to 5-(x+3)

OpenStudy (anonymous):

\[\large -2(x-3)\geq5-(x+3)\]?

OpenStudy (anonymous):

yes? no? sorta?

OpenStudy (anonymous):

i am going to take that as a "yes"

OpenStudy (anonymous):

multiply out first get \[-2x+6\geq 5-x-3\] combine like terms gives \[-2x+6\geq -x+2\] add \(2x\) to both sides \[6\geq x+2\] subtract \(2\) and end with \[6\geq x\]